A sports car moving at constant velocity travels 110 m in 5.1 s .

If it then brakes and comes to a stop in 4.5 s , what is the magnitude of its acceleration (assumed constant) in m/s2?

Express the answer in terms of "g's," where g=9.80m/s2.
Answer #1

Given: Distance = 110 m , Time = 5.1 s

Therefore, Initial Velocity = 110 m / 5.1 s = 21.57 m/s

Now we have ,

Initial velocity = 21.57 m/s , Time = 4.5 s , Final velocity = 0 m/s

So, |Acceleration| = | (0 m/s - 21.57 m/s)/4.5 s |

= 4.79 m/s2 = 0.49 g